# Strategic Analysis of Fair Rank-Minimizing Mechanisms with Agent Refusal Option

### v4

@@ 5 Concluding Remarks
We show that if an assignment is determined by the uniform rank-minimizing mechanism, which is a fair one, and agents can refuse the assignment and obtain the outside option instead, then the true preference is strategically dominated by an ODS. The adoption of such strategies by agents may result in inefficient assignments; that is, an agent who adopts an ODS refuses their assigned type, although some other agents desire that type.
We consider a public school choice system as an example. Our result implies that a student may have an incentive to reveal truly unacceptable schools as acceptable schools, and they may be assigned to one of them and then refuse the admission even if some other students desire the school. In such a case, the school may re-enroll students, but even in that case, it would cause market disruption.
@@
Therefore, if resolving the problem in that way is difficult, it is necessary to consider alternative mechanisms. First, as shown in this study, if we modify the uniform rank-minimizing mechanism, then truth-telling is not strategically dominated by any other strategies, including the outside option demotion ones. However, in this modification, it is necessary to ensure that increasing the number of acceptable types can result in a disadvantage. Thus, this modification may lead to the opposite problem, where agents strategically decrease the number of acceptable types. Second, we consider efficient and fair mechanisms other than the fair rank-minimizing mechanisms, such as the (general) probabilistic serial mechanism introduced by Bogomolnaia and Moulin (2001) and Budish et al. (2013), because it is fair and ordinally efficient, although this efficiency property is weaker than the rank-minimizing.
@@
$R\left(o,r_{a}\right)$ $<$ $R\left(\varnothing,r_{a}\right)<R\left(o^{\prime},r_{a}\right)$
$\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}$$\geq$ $\left|\mathcal{A}\right|,$ (9)
for some $o,o^{\prime}\in\mathcal{O}$. Fix $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ arbitrarily. Moreover, we let $d_{a}\in\mathcal{D}_{a}$ be such that $R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},d_{a}\right)$.
We show the following fact.
Claim 3
Let $y$ and $y^{\prime}$ be such that $y_{a\varnothing}=y_{ao^{\prime}}^{\prime}=1$ and $y_{a^{\prime}\bar{o}}=y_{a^{\prime}\bar{o}}^{\prime}$ for all $a^{\prime}\in\mathcal{A}$ and all $\bar{o}\in\mathcal{O}$. Then, $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ if and only if $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
Proof. By the construction,
$RV\left(y,\left(r_{a},p_{-a}\right)\right)=RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)\text{.}$
First, suppose $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. Then, by (9), $y^{\prime}$ is an allocation. Suppose not; that is, there is $y^{\prime\prime}\in\mathcal{Y}$ such that $RV\left(y^{\prime\prime},\left(d_{a},p_{-a}\right)\right)<RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)$. Then, by (9), $o^{\prime\prime}$ such that $y_{ao^{\prime\prime}}^{\prime\prime}=1$ satisfies $R\left(o,d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$. Let
$o^{\prime\prime\prime}=o^{\prime\prime}\text{ if }o\neq o^{\prime},\text{ }o^{\prime\prime\prime}=\varnothing\text{ if }o^{\prime\prime}=o^{\prime}\text{.}$
Moreover, let $y^{\prime\prime\prime}$ be such that $y_{ao^{\prime\prime\prime}}^{\prime\prime\prime}=1$ and $y_{a^{\prime}\bar{o}}^{\prime\prime\prime}=y_{a^{\prime}\bar{o}}^{\prime\prime}$ for all $a^{\prime}\in\mathcal{A}$ and all $\bar{o}\in\mathcal{O}$. Then,
$RV\left(y^{\prime\prime\prime},\left(r_{a},p_{-a}\right)\right)\leq RV\left(y^{\prime\prime},\left(d_{a},p_{-a}\right)\right)<RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)=RV\left(y,\left(r_{a},p_{-a}\right)\right)$
contradicting $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$.
Second, suppose $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$. Suppose not; that is, there is $y^{\prime\prime}\in\mathcal{Y}$ such that $RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right)$. Then, by (9), $o^{\prime\prime}\in\mathcal{O}$ such that $y_{ao^{\prime\prime}}^{\prime\prime}=1$ and $R\left(o^{\prime\prime},r_{a}\right)\leq R\left(\varnothing,r_{a}\right)$. Let
$o^{\prime\prime\prime}=o^{\prime\prime}\text{ if }o^{\prime\prime}\neq\varnothing,\text{ }o^{\prime\prime\prime}=\bar{o}\text{ if }o^{\prime\prime}=\varnothing\text{,}$
where $\bar{o}$ satisfies $\sum y_{a\bar{o}}^{\prime\prime}<q_{\bar{o}}$ and $R\left(\bar{o},d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$. There exist such an $\bar{o}$, because (9) is satisfied. Then, let $y^{\prime\prime\prime}$ be such that $y_{ao^{\prime\prime\prime}}^{\prime\prime\prime}=1$ and $y_{a^{\prime}\bar{o}}^{\prime\prime\prime}=y_{a^{\prime}\bar{o}}^{\prime\prime}$ for all $a^{\prime}\in\mathcal{A}$ and all $\bar{o}\in\mathcal{O}$. Then,
$RV\left(y^{\prime\prime\prime},\left(d_{a},p_{-a}\right)\right)\leq RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)=RV\left(y,\left(d_{a},p_{-a}\right)\right)$
contradicting $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
Therefore, $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ if and only if $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$. Q.E.D.
By Claims 1 and 3, we have Proposition 3.
Proof of Lemma 4
We assume that the assumptions of Lemma 4 is satisfied. First, by Proposition 3, for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and all $\bar{o}\in\mathcal{O}$,
$g\left(f^{U}\left(e_{a},p_{-a}\right),r\right)_{a\bar{o}}=g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)_{a\bar{o}}.$
Second, by Theorem 2, $r_{a}$ is strategically dominated by $d_{a}$ with the refusal option. Therefore, $e_{a}$ is also strategically dominated by $d_{a}$ with the refusal option.
Proof of Proposition 4

### v5

@@ 5 Concluding Remarks
We show that if an assignment is determined by the uniform rank-minimizing mechanism, which is a fair one, and agents can refuse the assignment and obtain the outside option instead, then truth-telling is strategically dominated by an ODS. The adoption of such strategies by agents may result in inefficient assignments; that is, an agent who adopts an ODS refuses their assigned type, although some other agents desire that type.
We consider a public school choice system as an example. Our result implies that a student may have an incentive to reveal truly unacceptable schools as acceptable schools, and they may be assigned to one of them and then refuse the admission even if some other students desire the school. In such a case, the school might re-enroll students, but even in that case, it would cause market disruption.
@@
Therefore, if resolving the problem in that way is difficult, it is necessary to consider alternative mechanisms. First, as shown in this study, if we modify the uniform rank-minimizing mechanism, then truth-telling is not strategically dominated by any other strategies, including the outside option demotion ones. However, in this modification, it is necessary to ensure that increasing the number of acceptable types can result in a disadvantage. Thus, this modification may lead to the opposite problem, where agents strategically decrease the number of acceptable types.
Second, we consider efficient and fair mechanisms other than the fair rank-minimizing mechanisms, such as the (general) probabilistic serial mechanism introduced by Bogomolnaia and Moulin (2001) and Budish et al. (2013), because it is fair and ordinally efficient, although this efficiency property is weaker than the rank-minimizing.
Finally, our study does not address reallocation after assignment refusals. However, Feigenbaum et al. (2020) and Afacan (2022) consider dynamic markets, such as school choice markets with a second-round admission process, where reallocation after refusals is possible. Incorporating such a framework could enhance the practical applicability of our findings. Specifically, analyzing how refused objects are redistributed and how agents’ strategic behavior changes in response could provide new insights into the design of rank-minimizing mechanisms.